OpenStudy (ally_b):

1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 = quantity n times quantity six n squared minus three n minus one all divided by two Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false.

OpenStudy (welshfella):

Lets assume its trus for n = k terms so Sk = k (6k - 3k - 1) ------------- 2 lets add the next term (k+1)th term S(k+1) = k(6k - 3k - 1) + 3((k + 1) - 2)^2 ----------- 2 Now we have to try and convert this to same form as the expression for Sk but with every k replaced by (k+1)

OpenStudy (welshfella):

* thats 6k^2 in the formula NOT 6k

OpenStudy (welshfella):

now we are aiming for Sk+1 = (k+1) (6(k+1)^2 - 3(k+1) - 1) -------------------------- 2 which is Sk with (k+1) replacing k. If we can convert the previous Sk+1 expression to this we have in the way to the proof.

OpenStudy (ally_b):

So the statement is true for all positive integers? @welshfella

OpenStudy (welshfella):

we need to do the algebra first to see if the 2 formula for S(k + 1) are the same . Its pretty messy I'm afraid.

OpenStudy (ally_b):

So what would I do next to solve the problem? @welshfella

OpenStudy (welshfella):

simplify the 2 expressions for S(k+1) and see if they are the same.

OpenStudy (ally_b):

okay thank you! @welshfella

OpenStudy (welshfella):

you'll find that they are the same so if formula is true for n=k then its true for n = k+1 And we know the formula is trus for k = 1 therfore its true for k = 2 . 3 and so on So ny induction the formula is true for all positive integers n